### v3

# A Database of Continued Fractions of Polynomial Type

Abstract
We describe a database of 1779 continued fractions with polynomial coefficients, of which more than 1500 are new, both for interesting constants and for transcendental functions, and provide the database inside the  source of the paper. Look in particular at the section “Table of Contents”.
1 Introduction
Please download the  source of this paper since it contains the database described here, otherwise this paper is useless.
This is not quite a standard mathematical paper, but I believe it will be useful (and hopefully become a standard reference) to any mathematician interested in continued fractions, in particular attached to interesting constants or functions.
I have developed an extensive Pari/GP package for dealing with continued fractions (abbreviated CF in the sequel), for now accessible only as a GIT branch under the name henri-ellCF (please read the Pari/GP installation documentation to obtain this). Thanks to this package, I have done several things.
  1. 1.
    First, starting from the 250 or so CFs with polynomial coefficients (see below for the precise definition) existing in the literature, I have succeeded in creating more than 1500 additional ones, some of them I believe quite interesting.
  2. 2.
    Second, for almost every CF in the corresponding database, I have computed the precise rate of convergence of the partial quotients. This is almost never done in the literature, although many texts deal with convergence of CFs. To give an almost random example, we have
    $\log(2)=3/(5-9\cdot 1^{2}/(15-9\cdot 2^{2}/(25-9\cdot 3^{2}/(35-9\cdot 4^{2}-\cdots))))\;,$
    and in addition to this CF the database contains the following:
    $\log(2)-\dfrac{p(n)}{q(n)}=\dfrac{\pi/3}{9^{n}}\left(1-\dfrac{5/16}{n}+\dfrac{105/512}{n^{2}}-\dfrac{1725/8192}{n^{3}}+\cdots\right)\;.$
  3. 3.
    Third, in many cases, I have given some explanation on the origin of the CF, in particular when it has been obtained by Apéry acceleration techniques, see [2].
  4. 4.
    Fourth, in many cases the CF can also be expressed as a hypergeometric series, which is given.
I have created a file containing these 1779 continued fractions together with their relevant information, which is readable from Pari/GP with the command V=readvec("file.gp"), if file.gp is the name given to the file (it is of course also readable by other programs). To obtain this file, from the arXiv repository from which you downloaded the pdf you must also obtain the  source, and the file is entirely contained between the last two comment commands.
2 The Database in Detail
I will now describe the precise format of the database of continued fractions. It is intended to be used by Pari/GP, but of course since the format is transparent it can be used with any other software through an appropriate interface.
2.1 Computer Representation of Continued Fractions
Let $(a(n))_{n\geq 0}$ and $(b(n))_{n\geq 0}$ be two sequences of complex numbers. The continued fraction $(a,b)$ is the formal expression
$(a,b)=a(0)+\dfrac{b(0)}{a(1)+\dfrac{b(1)}{a(2)+\dfrac{b(2)}{a(3)+\ddots}}}$
written more compactly as $(a,b)=a(0)+b(0)/(a(1)+b(1)/(a(2)+\cdots))$. Note that this notation is different from that used by other authors such as [4]. I do not claim that my notation is better but simply that it is the one used in the present paper. We will always assume that $b(n)\neq 0$ for all $n$ (otherwise the CF terminates), and that $a(n)=0$ only for a finite number of $n$. We write as usual formally $p(n)/q(n)=a(0)+b(0)/(a(1)+b(1)/(a(2)+\cdots+b(n-1)/a(n)))$.
Definition 2.1
A CF $((a(n))_{n\geq 0},(b(n))_{n\geq 0})$ is said to have polynomial coefficients if there exists a period $T$ such that for every $j$ with $0\leq j<T$, for $n$ sufficiently large both $a(nT+j)$ and $b(nT+j)$ are rational functions of $n$ with rational coefficients.
The name polynomial is justified since it is clear that any CF with coefficients in $\mathbb{Q}(n)$ is equivalent (i.e., a CF with the same $p(n)/q(n)$) to one with coefficients in $\mathbb{Z}[n]$.
In practice $T$ will be equal to $1$ or $2$, so for $n$ sufficiently large either $a(n)$ and $b(n)$ are rational functions, or $a(2n)$, $a(2n+1)$, $b(2n)$, and $b(2n+1)$ are rational functions.
It is evidently easy to contract a CF of period 2 to a CF of period 1, but the point of keeping a period 2 CF is that often its coefficients are simple, while those of the contracted CF are complicated.
Note that (of course excluding finite and periodic CFs) most explicit CFs occurring in the literature have polynomial coefficients, with the notable exception of $q$-continued fractions such as Ramanujan’s
$R(q)=q^{1/5}/(1+q/(1+q^{2}/(1+q^{3}/(1+\cdots))))\;,$
which are fascinating but will not be considered here.
Let $(a,b)=((a(n))_{n\geq 0},(b(n))_{n\geq 0})$ be a continued fraction with polynomial coefficients. If $a$ (similarly $b$) has period $1$, i.e., if $a(n)$ is a polynomial for $n\geq n_{0}$ for some $n_{0}$, we will represent $a$ as the vector
$A=[a(0),a(1),\dotsc,a(n_{0}-1),a(n)]\;,$
where for $i\leq n_{0}-1$ the $a(i)$ do not depend on $n$, while $a(n)$ is a polynomial (exceptionally also a rational function) in the reserved variable $n$. Note that the variable $n$ is reserved for this purpose and cannot be used elsewhere. It would perhaps have been more logical to use the variable $x$ instead, but the variable $n$ is almost universally used in texts so I have chosen to use it.
Note also that the representation is not unique, for instance we also have $A=[a(0),a(1),\dotsc,a(n_{0}-1),a(n_{0}),a(n)]$.
For instance, the continued fraction
$\log(2)=3/(5-9\cdot 1^{2}/(15-9\cdot 2^{2}/(25-9\cdot 3^{2}/(35-9\cdot 4^{2}-\cdots))))$
given above is represented by
[[0,10*n-5],[3,-9*nˆ2]]
If $A$ has period $2$, i.e., if $a(2n)$ and $a(2n+1)$ are possibly different polynomials for $n$ large, we will represent $A$ as a vector of $2$-component vectors
$A=[[a(0),a(1)],\dotsc,[a(2m_{0}-2),a(2m_{0}-1)],[a(2n),a(2n+1)]]\;.$
For instance, for the continued fraction
$\pi=4-4/(3+9/(5+4/(7+25/(9+16/(11+49/(13+36/(15+\cdots)))))))\;,$
AA has period 1 but $B$ has period 2, so we can represent the CF by
[[4,2*n+1],[[-4,9],[(2*n)ˆ2,(2*n+3)ˆ2]]]
Note that in this case all the components of $B$ must be two-component vectors. For instance the above could not be written as
[[4,2*n+1],[-4,[(2*n+1)ˆ2,(2*n)ˆ2]]]
2.2 The Database Formats
First, a warning. We reserve the following variable names which must never be used in working with the CFs of this file:
  1. 1.
    “n”, which is universally used to denote the CF index.
  2. 2.
    “d”, which denotes the radicand of a quadratic extension, see below the explanation for v[3][1].
  3. 3.
    “x” which stands for $1/n$ and “y” which stands for $1/n^{1/2}$ in the asymptotic expansions $A$, see the explanation for v[3][4].
There are three kinds of inputs in the file. By far the most important ones are of course the CFs themselves, which are also the most numerous.
$\bullet$ Format of the Continued Fractions Entries
Such an entry, let’s call it v, has eight components, which we now describe. To make the explanation clearer, we use the following example:
[(z)->cosh(Pi*z),[[1,1/2,4*n^2-8*n+(2*z^2+5)],
[2*z^2,-4*n^4+8*n^3+(-4*z^2-6)*n^2+(4*z^2+2)*n+(-z^2-1/4)]],
[1,[0,1,0,1],(z)->z^2*cosh(Pi*z),
1-1/2*z^2*x+(1/6*z^4-1/6*z^2-1/12)*x^2+(-1/24*z^4+1/6*z^2+1/12)*x^3+O(x^4)],
[[1,4*z^2,[1/2-I*z,1;1/2+I*z,1;3/2,-2],1],
 [1,-4*z^2,[-1/2,2;1/2-I*z,-1;1/2+I*z,-1],1]],
[],"3.2.4",["AP->3.2.5","APD->3.2.4.5"],""];
  1. 1.
    v[1] is a closure giving the limit of the CF. Note that in CFs involving variables, the CF may converge to that limit only in a certain range of the variables, but we have not given that (we may in a future version).
  2. 2.
    v[2] is the $[a(n),b(n)]$ of the CF in the format explained above, with one important difference: consider for instance the entry 1.5.10 beginning with:
      [k->zeta(k),[0,n^k+(n-1)^k],[1,-n^(2*k)]]
    Giving this to GP would make it choke (give an error message) because nˆk is not a valid GP input when k is a variable. Thus we replace it in the file by
      [k->zeta(k),[k->[0,n^k+(n-1)^k],k->[1,-n^(2*k)]]]
    which is a valid input since the closure is not evaluated.
  3. 3.
    v[3] is itself a 4-component vector [d,[F,E,D,P],C,A] giving the exact speed of convergence of the CF, including a few terms of its asymptotic expansion. If $L$ denotes the value of the CF and $p(n)/q(n)$ are the partial quotients, this notation means that
    $L-\dfrac{p(n)}{q(n)}=\dfrac{C}{n!^{F}E^{n}e^{\sqrt{Dn}}n^{P}}A\;,$
    with the following conventions:
    1. (a)
      v[3][1] is a number or a polynomial (very often 1) with the following meaning: often, the speed of convergence and the asymptotic expansions involve elements of a quadratic extension (of $\mathbb{Q}$ or of $\mathbb{Q}[z]$ for instance). To avoid using square roots, we set v[3][1] to the radicand, and we use the specific letter d to indicate the square root of v[3][1]. Thus, we must avoid the letter d in variables of CFs. In the running example, we have v[3][1]=1 so no quadratic extension.
    2. (b)
      v[3][2] gives the speed of convergence in the above FEDP format, using d if necessary. The running example has v[3][2]=[0,1,0,1], meaning that we are in case $P^{+}$ with $P=1$. In some rare cases, the speed does not make any sense, in which case we set v[3][2]=0. In some other rare cases, the CF does not converge for real values of the variables (which is implicitly what is assumed), in which case we set v[3][2]=[0,0,0,0].
    3. (c)
      v[3][3] gives the constant $C$ such that $S-p(n)/q(n)\sim C/[FEDP]$, or 0 if unknown or has not been computed. Note that $C$ is always given as a closure, first because in some cases (as for the CF itself) it is not possible to do otherwise in GP, and second because writing C=()->Pi is both much clearer to read than C=3.1415... and also allows to have C to any accuracy.
    4. (d)
      v[3][4] gives the asympotic expansion A=1+..., such that Sp(n)/q(n)=(C/[FEDP])AS-p(n)/q(n)=(C/[FEDP])A, or 0 if unknown or not computed. The reserved letter x means $1/n$ and the reserved letter y means $1/n^{1/2}$.
      Thus, the above example says that
      $\cosh(\pi z)-\dfrac{p(n)}{q(n)}=\dfrac{z^{2}\cosh(\pi z)}{n}\left(1-\dfrac{z^{2}/2}{n}+\dfrac{z^{4}/6-z^{2}/2-1/12}{n^{2}}+\cdots\right)$
  4. 4.
    In a large number of cases, the CF or its inverse corresponds to a more or less explicit hypergeometric series S given in a special format explained below. Thus, v[4] is set to 0 if neither CF nor its inverse has such a series, to S if only the CF has a series, to the two-component vector [S,T] if both have a series (with T the series for the inverse), and finally the two-component vector [[],T] if only the inverse has a series.
    These series are given in the following format: they are all of the form
    $S=u+\sum_{n\geq 0}h(n)\prod_{1\leq i\leq g}(a_{i})_{n}^{e_{i}}w^{n}\;,$
    where $h$ is a rational function, $e_{i}\in\mathbb{Z}$, and as usual $(a_{i})_{n}$ denotes the rising Pochhammer symbol $\Gamma(a_{i}+n)/\Gamma(a_{i})$ (note that the series always starts at $n=0$). This is represented by
    S=[u,h(n),[a_1,e_1;a_2,e_2;...a_g,e_g],w]
    (when $g=0$, the empty matrix is simply set to $1$). Note that for convergence, we always have $\sum_{1\leq i\leq g}e_{i}\leq 0$. Thus, the above example says that
    $\cosh(\pi z)$ $=1+4z^{2}\sum_{n\geq 0}\dfrac{(1/2-iz)_{n}(1/2+iz)_{n}}{(3/2)_{n}^{2}}\text{\quad and}$
    $\dfrac{1}{\cosh(\pi z)}$ $=1-4z^{2}\sum_{n\geq 1}\dfrac{(-1/2)_{n}^{2}}{(1/2-iz)_{n}(1/2+iz)_{n}}$
    The fact that the series begins at $n=0$ has been chosen for uniformity, but often leads to slightly more cumbersome formulas, for instance instead of $\zeta(2)=\sum_{n\geq 1}1/n^{2}$ we write instead $\zeta(2)=\sum_{n\geq 0}1/(n+1)^{2}$, hence represented by [0,1/(n+1)ˆ2,1,1].
  5. 5.
    v[5] is used for testing. When empty, it means that the default values for the variables used in the CF can be used (presently $2/3$, $3/4$, $1/2$, 22 for closures with arity less or equal to $4$). If it is a vector of values, these are taken instead of the default values. For instance for a closure of arity $2$, v[5]=[2/3,-1] means that one can take the default value $2/3$ for the first variable, but probably cannot use $3/4$ for the second, but 1-1 instead. Other values of v[5] have specific meanings which may change:
    • v[5]=1: Bernoulli type sequence, test expansion.
    • v[5]=2: Jacobi type sequence, only test expansion.
    • v[5]=3: Jacobi type sequence, test expansion and numerics, (2,1/2).
    • v[5]=4: Change z into 1/z, multiply by z, test expansion.
    • v[5]=5: Choose $z=2i$, multiply by i-i, test.
    Note that in the present version, v[5]=4 and v[5]=5 occur each for a single CF.
  6. 6.
    v[6] is the  label used for the CF, so as to be able to find it easily. It is very useful to write a trivial little script to find conversely the given entry of the big vector V corresponding to a given label. Note that the labels are rather arbitrary (but of course unique) since they were given after discovery of new CFs, and more or less in increasing order.
  7. 7.
    v[7] is a 2-component vector [A1,AD1] whose entries are strings describing the behavior of the CF under Apéry acceleration: A1 is for the Apéry diagonal process, and AD1 for the Apéry dual. The meanings of the strings are as follows: “NO” if Apéry is not applicable, “BAcomp” if it is but the Bauer–Muir method introduces complicated denominators, “AP→label” (“APSI→label” if followed by a simplification) and “APD→label” if Apéry leads to the corresponding labeled CF, (“APD→SELF” it is self-dual), “APcomp” and “APDcomp” if the result is too complicated to be included in the encyclopedia so with no label, and a possible additional “sim” if similar to.
  8. 8.
    v[8] is a (usually empty) string giving any additional information.
$\bullet$ Format of Parametrized Continued Fractions
Many CFs are part of families of inequivalent CFs depending on one or more parameters. These families are included in the file in a different format, since the initial values of $a(n)$ and $b(n)$ are not given. We have only given such families for CFs of period 1.
The corresponding vector v has now only five components, and again to make the explanation clearer we give an example:
[[()->Pi,2*k+2,(2*n-1)*(2*n+2*k-4*u-1)]
[1,[0,-1,0,k+1]],"1.2.1",[],"u<=k/2"];
  1. 1.
    The first component v[1] is a three-component vector: v[1][1] is the closure giving the limit $S$, and v[1][2] and v[1][3] are polynomials giving $a(n)$ and $b(n)$ for $n$ sufficiently large. This means that for suitable initial values $a(0),a(1),\dotsc$, and $b(0),b(1),\dotsc$, and for given nonnegative integers $k$ and $u$, the corresponding CF converges to $S$. Equivalently, the CF [[v[1][2]],v[1][3]] converges to a limit of the form $(aS+b)/(cS+d)$ with $a$, $b$, $c$, $d$ integral and $ad-bc\neq 0$. For instance, choosing $u=0$ and $k=0,1,2$, or $u=1$ and $k=2$, the above parametrized CF says that the limits of the CFs
    $2+1/(2+1/(2+9/(2+25/(2+49/(2+81/(2+\cdots))))))\;,$
    $4-1/(4+3/(4+15/(4+35/(4+63/(4+99/(4+\cdots))))))\;,$
    $6-3/(6+5/(6+21/(6+45/(6+77/(6+117/(6+\cdots))))))\;,\text{\quad and}$
    $6+1/(6+1/(6+9/(6+25/(6+49/(6+81/(6+\cdots))))))$
    are all of the form $(A\pi+B)/(C\pi+D)$ with $AD-BC\neq 0$ and $A$, $B$, $C$, and $D$ in $\mathbb{Z}$.
  2. 2.
    v[2] contains the information on the speed of convergence in the shortened format [d,[F,E,D,P]], since the constant C and the asymptotic expansion A vary in a parametric family so are not given.
  3. 3.
    v[3] is the label where the parametric CF occurs (same as v[5] for ordinary CFs).
  4. 4.
    v[4] is used for testing, but for now is the empty vector.
  5. 5.
    v[5] is a string giving some additional information.
Note that for integer $k$, the CF $(a(n+k),b(n+k))$ is essentially the same as the CF $(a(n),b(n))$, so is of course not considered as a new CF, so this trivial parametrization is not included.
$\bullet$ Format of Definitions
In some of the CFs, it is necessary to add some function definitions, either because the corresponding functions do not exist in Pari/GP (for example bernfrac(k) and eulerfrac(k) exist, but not tanfrac(k) for the tangent numbers), or simply because the existing expression is cumbersome. The command V=readvec("file.gp"); will have two effects: first, it defines the function so that it can be used in the CFs, and second it is kept as an element of V as a closure (but without its name). Thus, if you do not know the meaning of a function, typing ?functionname will give you an explanation or a formula for it.
More Details
Here is in more detail the different types of vectors $v$ that we can encounter, with examples.
  1. 1.
    Definitions: as explained above. These are closures. Example:
      R1(k,n)=(n^k*(2*n+1)+(n-1)^k*(2*n-3))/(2*n-1);
  2. 2.
    Parametrized continued fractions: as explained above. These are vectors with five components. Example:
    [[()->2^(1/3),7*n-5+k+v+4*u,-4*(n+u)*(3*n+v-2)],
                [1,[0,4/3,0,u+(5-v)/3+2*k]],"1.1.1",[],""];
  3. 3.
    Ordinary CF with $a$ or $b$ expressed as a closure: These are vectors $v$ with eight components such that $v[2][1]$ or $v[2][2]$ is a closure. Example:
    [k->zeta(k),[k->[0,n^k+(n-1)^k],k->[1,-n^(2*k)]],[1,[0,1,0,k-1],
    k->1/(k-1),1-((k-1)/2)*x+sum(m=1,4,binomial(k+2*m-2,2*m)*
    bernfrac(2*m)*x^(2*m),O(x^10))],0,[3],"1.5.10",["NO","NO"],""];
  4. 4.
    Ordinary CF with the last component of $a$ and $b$ a closure. Example:
      [()->1,[[0,n->f(n)],[n->f(n+1)+1]],[0,0,0,0],
                                  0,[],"2.1.0.1",["NO","NO"],""];
  5. 5.
    Ordinary CF of period 2 or more: These are vectors $v$ with eight components such that the first (hence any) component of $v[2][1]$ or $v[2][2]$ is a vector. Example:
    [()->2^(1/3),[[[0,5],[2*n,12*n+6]],[[6,-1/3],[-(n+1)*(3*n+2),
    -n*(3*n+1)]]],[2,[0,(1+d)^2,0,0],()->2^(1/3)*3^(1/2)/(1+sqrt(2))^2,
    1+(9*d/8)*x+(-9*d/8+81/64)*x^2+(67105*d/41472-81/32)*x^3+O(x^4)],
    0,[],"1.1.6",["NO","NO"],""];
  6. 6.
    Ordinary CF of period 1 with no closures. Example:
    [()->3^(1/2),[[3/2,(2*n+1)^2],[2,-n^2*(n+2)^2]],[3,[0,(2+d)^2,0,0],
    ()->2*sqrt(3)/(2+sqrt(3))^3,1+d*x+(3/2-3*d/2)*x^2+(3*d-9/2)*x^3
    +(45/4-27*d/4)*x^4+O(x^5)],0,[],"1.1.0.5",
    ["NO","NO"],"Infinitely contractible"];
    A second example with Apéry information and series for the inverse:
    [()->2^(1/3),[[1/2,7*n-5],[1,-12*n^2+8*n]],[1,[0,4/3,0,5/3],
    ()->2^(8/3)/gamma(1/3),1-55/9*x+170/3*x^2-1579490/2187*x^3
    +230596751/19683*x^4+O(x^5)],[[],[0,2,[-2/3,1;1,-1],3/4]],
    [],"1.1.1",["APSI->1.1.8","APD->SELF"],""];
    where the string “APSI→1.1.8” indicates that after simplification, Apéry leads to label 1.1.8, and “APD→SELF” that it is self-dual for Apéry. The series entry says that $2^{-1/3}=2\sum_{n\geq 0}((-2/3)_{n}/n!)(3/4)^{n}$, which is simply the special case $x=3/4$ of the expansion of $2(1-x)^{2/3}$.
Here is a little script which outputs a number from 1 to 6 giving the type of a vector $v$:
/* Find the type number of an entry in the file */
typevec(v)=
{ my(tv=type(v),lv=#v,a,b);
  if(tv=="t_CLOSURE",return(1));
  if(tv!="t_VEC"||(lv!=5&&lv!=8),error("incorrect entry"));
  if(lv==5,return(2));
  [a,b]=v[2];
  if(type(a)=="t_CLOSURE"||type(b)=="t_CLOSURE",return(3));
  if(type(a)!="t_VEC"||type(b)!="t_VEC",error("incorrect entry 2"));
  if(type(a[#a])=="t_CLOSURE"||type(b[#b])=="t_CLOSURE",return(4));
  if(type(a[1])=="t_VEC"||type(b[1])=="t_VEC",return(5));
  return(6); }
2.3 Table of Contents
The complete GP file of the 1779 CFs and definitions together with convergence and additional data can be obtained by obtaining the  source and saving the file contained between the comment statements which are at the end of the source file, and then reading it under GP with the command V=readvec("file.gp") if file.gp is the name that you gave to your saved file. The source file itself is divided into sections (included as comments so not read by GP) which helps to order the different CFs. The sections are the following, together with the range of corresponding indices of the vector V:
Definitions: 1–45
Constants: $2^{1/3}$: 46–82
Constants: $\pi$, $\log(2)$, and Periods of Degree 1: 83–217
Constants: $\pi^{2}$, $G$, and Periods of Degree 2: 218–316
Constants: $\pi^{3}$, $\zeta(3)$, and Periods of Degree 3: 317–345
Constants: $\pi^{4}$: 346–352
Constants: $\zeta(k)$ and $\pi^{k}$ for $k\geq 5$: 353–356
Constants: Linear Combinations of Zeta- and $L$-Values: 357–633
Miscellaneous Constants Related to $L$-Values: 634–636
Constants Related to Powers of $e=\exp(1)$: 637–674
Constants Involving Gamma Quotients: 675–999
Constants Coming from Bessel Functions: 1000–1019
Trivial Functions with an Arbitrary Sequence: 1020–1030
Other Trivial Functions: 1031–1043
Exponential Functions: 1044–1071
Logarithm Functions: 1072–1096
Power Functions: 1097–1121
Direct Hyperbolic and Trigonometric Functions: 1122–1191
Inverse Hyperbolic and Trigonometric Functions: 1192–1220
Gamma Function: One Variable: 1221–1275
Gamma Function: Several Variables: 1276–1291
Function $\psi(z)=\Gamma^{\prime}(z)/\Gamma(z)$: 1292–1324
Function $\beta(z)=(\psi((1+z)/2)-\psi(z/2))/2$: 1325–1400
Function $\psi^{\prime}(z)$: 1401–1426
Function $\beta_{1}(z)=-\beta^{\prime}(z)=\psi^{\prime}(z/2)/2-\psi^{\prime}(z)$: 1427–1437
Function $\psi^{\prime\prime}(z)$: 1438–1455
Function $\beta_{2}(z)=-\beta^{\prime}_{1}(z)=\psi^{\prime\prime}(z)-\psi^{\prime\prime}(z/2)/4$: 1456–1457
Function $\psi^{\prime\prime\prime}(z)$: 1458
Ordinary Generating Functions for Bernoulli Type Sequences: 1459–1583
Incomplete Gamma Functions: 1584–1595
Error Function and Related: 1596–1612
Jacobi Elliptic Functions: Laplace Transforms: 1613–1648
Complete Elliptic Integrals: 1649–1665
More Integrals: 1666–1688
Bessel and Related Functions: 1689–1714
General Hypergeometric Functions: $_{0}F_{1}$: 1715–1718
General Hypergeometric Functions: $_{1}F_{1}$: 1719–1726
General Hypergeometric Functions: $_{2}F_{0}$: 1727–1740
General Hypergeometric Functions: $U$: 1741–1742
General Hypergeometric Functions: $_{2}F_{1}$: 1743–1760
General Hypergeometric Functions: $_{2}F_{1}$ with $b=1$: 1761–1775
General Hypergeometric Functions: $_{3}F_{2}$ with $z=1$: 1776–1779
Note that if you do not understand the name of a function, you can use the GP command ?name to obtain an explanation or a formula. For instance, typing ?tanfrac returns
tanfrac =
  (k)->if(k==0,0,2^(k+1)*(2^(k+1)-1)*bernfrac(k+1)/(k+1))
Warnings. Evidently in such a large database essentially put together “by hand” it is unavoidable that there are typographical errors. But apart from this, two warnings are necessary.
  1. 1.
    First, the speed of convergence has been obtained using heuristic methods, although I believe that the results are correct (in particular, for CFs for functions, the indicated speed of convergence often depends on the range of the variables). In addition, the constant $C$ has been obtained using linear dependence algorithms (typically the lindep function of Pari/GP): it should be correct, but I have not tried to prove anything.
  2. 2.
    Second, many of the new CFs have been obtained rigorously from existing ones, for instance by Bauer–Muir–Apéry type methods, but many others have been obtained “à la Ramanujan”, in particular not rigorously, although I am quite confident of their correctness. Thus it is absolutely impossible for me to give the origin, and a fortiori the proofs, of the CFs in the database.
Evidently I would more than welcome corrections and/or additions. Since each CF has a label number, please refer to it when sending comments. Ideally, this file should be put on a web site which could be regularly updated by users, but I do not have the necessary knowledge to do this.
2.4 Beautifying the Continued Fractions
From a given CF in the database format, it is not difficult to create a “beautified” version which is more human-readable, and I advise the reader to do so. In fact, the author has done so and added a large quantity of additional information, thus creating a 630 page book which may appear someday. In particular, some CFs may not be understandable without this additional information, but I believe that the present file is sufficiently useful as is.
Here is an example of beautification. Entries V[1051] and V[1052] are (newlines inserted for better readability):
[(z)->exp(Pi*z/2),[[1,-z+1,2],[2*z,4*n^2-4*n+(z^2+1)]],
[1,[0,-1,0,1],(z)->z*exp(Pi*z/2)*cosh(Pi*z/2),
1+(-1/16*z^2-1/4)*x^2+(1/128*z^4+7/64*z^2+5/16)*x^4+O(x^6)],
0,[],"3.1.5.0",["AP->3.1.5.2","APD->3.1.5.1"],""];
[[()->exp(Pi*z/2),4*k+2,4*n^2-4*n+(z^2+1)],
[1,[0,-1,0,2*k+1]],"3.1.5.0",[],""];
A possible beautified version is:
3.1.5.0
[(z)->exp(Pi*z/2),[1,1-z,2],[2*z,z^2+(2*n-1)^2]]
$e^{\pi z/2}=1+\dfrac{2z}{-z+1+\dfrac{z^{2}+1}{2+\dfrac{z^{2}+9}{2+\dfrac{z^{2}+25}{2+\dfrac{z^{2}+49}{2+\dfrac{z^{2}+81}{2+\ddots}}}}}}$
Convergence type $P^{-}$ with $P=1$ and $C=ze^{\pi z/2}\cosh(\pi z/2)$, so that
$e^{\pi z/2}-\dfrac{p(n)}{q(n)}\sim(-1)^{n}\dfrac{ze^{\pi z/2}\cosh(\pi z/2)}{n}\;.$
$A=1+(-z^{2}/16-1/4)/n^{2}+(z^{4}/128+7z^{2}/64+5/16)/n^{4}+\cdots$
Parametric family for $k\geq 0$:
[(z)->exp(Pi*z/2),4*k+2,z^2+(2*n-1)^2]
Convergence type $P^{-}$ with $P=2k+1$.
Apéry accelerates to 3.1.5.2, Apéry dual is 3.1.5.1.
3 Some Remarks and Questions on the Database
In compiling and computing this database of continued fractions, a number of sporadic questions have arisen. Here are a few:
  1. 1.
    The CF given as an example for $3^{1/2}$ (label 1.1.0.5) has the curious property of being infinitely contractible: usually, computing the contraction of a CF of bidegree $(d_{1},d_{2})$ gives a CF of larger bidegree. For the present CF of bidegree $(2,4)$, the contractions become considerably more complicated, but can always be simplified (i.e., to a CF with the same partial quotients) to bidegree $(2,4)$. First, I do not know how to prove this, and second, I do not know how to characterize such CFs.
  2. 2.
    The parametric families, especially those for periods of degree $2$ and $3$ such as $\zeta(2)$ and $\zeta(3)$, have been obtained by searches. In many cases they are probably not complete, but in some cases there are so many that it would be desirable to classify them, which I have not always succeeded in doing. See [1] for detailed discussions of this.
  3. 3.
    In the case of logarithmic convergence, I do not know of any reasonable method to compute the constant $C$ such that $S-p(n)/q(n)\sim C/\log(n)$, except when the CF is completely explicit, or when it is the specialization of a CF with more variables (for instance, this is how I found that the constant $C$ for 1.3.22 is equal to $\pi^{3}/8$ and that of 1.4.2.4 is $\pi^{4}/28$).
  4. 4.
    When the convergence is precisely $S-p(n)/q(n)\sim C/n^{k}$ with integral $k$, I do not know of a systematic way to compute the asymptotic expansion $A$ to more than k1k-1 terms, except when the CF comes from a sum of explicit rational functions. In many cases, this is due to the presence of more complicated terms in the asymptotic expansion, typically involving $\log(n)$. In all other convergence cases (except of course the logarithmic case), including $C/n^{k}$ when $k\notin\mathbb{Z}$, there is an algorithmic way to compute $A$.
  5. 5.
    Since there are so many CFs for the functions $\beta(z)$ and $\beta_{1}(z)$, it is slightly surprising that there are no interesting ones for $\beta_{2}(z)=-\beta_{1}^{\prime}(z)$; nonetheless, I have given a couple of (uninteresting) ones.
  6. 6.
    Although $I_{\nu}(z)$ is an even function of $z$ (up to a power of $z$), the CF 5.1.7 does not seem to define an even function. This implies that there does not seem to exist a corresponding CF for $J_{\nu}(z)$ (not using complex numbers), and consequently that there does not seem to exist analogs for $J$ of the CFs 5.1.7.3 and 5.1.7.6 for $I$.
  7. 7.
    There exist many CFs involving the Bessel functions $J_{\nu}(z)$, $I_{\nu}(z)$, and $K_{\nu}(z)$, as well as $H_{\nu}(z)=J_{\nu}(z)\pm iY_{\nu}(z)$. Are there interesting ones involving $Y_{\nu}(z)$ alone ?
  8. 8.
    Using Apéry, it is not difficult to find a parametric family of CFs for $\cosh(\pi z/3)$, see 3.2.9. On the other hand, the very similar CF 3.2.10 for $\sinh(\pi z/3)/(z\sqrt{3}/2)$ does not accelerate nicely, so I have been unable to find a corresponding parametric family.
  9. 9.
    The CFs 4.1.14.A0 for $\Gamma(1/3)^{2}/\Gamma(2/3)$ and 4.1.14.C for $\Gamma(2/3)^{2}/\Gamma(1/3)$ are extremely similar and have the same origin. They can both be Apéry accelerated giving complicated formulas, but the former has a very simple Apéry dual 4.1.14.B, while the latter has a very complicated Apéry dual which therefore we have not given. Is there an analogue of 4.1.14.B for $\Gamma(2/3)^{2}/\Gamma(1/3)$ ?
  10. 10.
    Similarly, there is a beautiful rapidly convergent CF 4.1.14.C6 as well as a very simple hypergeometric series for $\Gamma(2/3)^{2}/\Gamma(1/3)$, but apparently nothing of the sort for $\Gamma(1/3)^{2}/\Gamma(2/3)$.